a student figures that he has a 26% chance of being let out of class late.if he leaves class late, there is a 82% chance thst he will miss his train. what is the probability that he gets out of class late and misses the train? the probabililty is
Answer #1

Let L is event that student is being let out of class late and T shows the event that student miss the train. So

P(L) = 0.26

P(T|L) = 0.82

So the probability that he gets out of class late and misses the train is

P(T and L) = P(T|L) P(L) = 0.82 * 0.26= 0.2132

Answer: 0.2132